How many liters of water must be added to 10L of 50% acid solution to produce 80% acid solution?
Answer by richwmiller(17219)

.5*10+0*x=.8*(x+10)
5=.8x+8
-3=.8x
multiply by ten to get rid of the decimal
-30=8x
-30/8=x
-3 6/8=x
-3.75 =x
3.75 liters of water need to be removed
This comes out with a negative amount so we must take away water.
You can't make a solution stronger (from 50% to 80%) by adding water.
But we can't take away just water , can we, from a solution so this problem can't be solved.
Maybe it can be done by distilling.